An LCD screen is purchased to be used in the manufacturing of a digital watch. The Standard Price​ (SP) for the LCD screen used is

$30.00.

During the month of​February,

2,600

screens were purchased and used. The materials price variance was

$6,500

unfavorable. The standard number of screens allowed for the actual number of watches manufactured during the period was

23,000

screens. The actual purchase price of each LCD screen may be closest to

A.

$32.50

per LCD screen.

B.

$27.50

per LCD screen.

C.

$3.54

per LCD screen.

D.

$30.40

per LCD screen.
Answer #1

Correct answer----(A) $ 32.50 per LCD screen

Material Price Variance

(

Standard Rate

-

Actual Rate

)

x

Actual Quantity

(

$          30.00

-

$         32.50

)

x

2600

-6500

Variance

$   6,500.00

Unfavourable-U

Variance = (SP-AP) x Actual quantity

6500=(30-AP) x 2600

6500= (78000 –AP2600)

AP2600=78000+6500

AP= 84500/2600

AP= $32.50